OpenStudy (anonymous):

28-(x+6)=8x-5 solve for x

OpenStudy (cristal):

27=9x divide by 9. x=3

OpenStudy (anonymous):

solution: 28-x-6=8x-5 -x-8x=-5+6-28 -7x=-39 x=39/7

jhonyy9 (jhonyy9):

the first is right x=3